# Elasticity without a reference state: continuum mechanics of active tension nets

### v2

@@ Results
subject to the BC $\operatorname{Re}[(\partial_{\xi}z)(\bar{\partial}_{\xi}\dot{z})]|_{\partial\Omega}=\partial_{t}\lambda_{g}|_{\partial\Omega}$. In other words, the anisotropic deformation of the metric $\partial_{t}\mu_{g}$ must be accommodated by a corresponding shear deformation. Tissue flow corresponds to a sequence of infinitesimal QC deformations – a Beltrami flow. The Beltrami flow equation is linear and can be solved by the method of Green’s functions, even for large deformations [26] (Appendix E).
We can restate the flow Eq. (40), formulated in Lagrangian $\xi$-coordinates, in Eulerian $z$-coordinates. Crucially, this allows computing the infinitesimal change in the embedding fully in terms of the present-time reference. We therefore reexpress the flow in the current $z(t)$ coordinates, so that $\dot{z}(z(t),\overline{z(t)},t)$ is the Eulerian velocity field. For brevity, we write $z=z(t)$. Eq. (40) becomes
$\bar{\partial}_{z}\dot{z}(z,\bar{z})=\frac{1}{2\lambda_{g}^{2}}\partial_{t}g_{zz}(z,\bar{z},t)$ (41)
(In the partial time-derivative $\partial_{t}$, the coordinates $z=z(t)$ are kept fixed. By definition of the isothermal $z(t)$, the total time derivative vanishes $\tfrac{d}{dt}g_{zz}(z(t),\bar{z}(t),t)=0$, but $\partial_{t}g_{zz}(z,\bar{z},t)$ does not.)
Eq. (41) can be brought into an alternate form which sheds light on the nature of the adiabatic flow. First expand Eq. (4) around $g_{z\bar{z}}=2\lambda_{g}^{2},\;g_{zz}=0$:
$\sigma_{\bar{z}z}$ $=p_{0}(2\lambda_{g}+(\partial_{t}g_{\bar{z}z})dt)$ (42a)
$\sigma_{zz}$ $=-\frac{p_{0}}{2\lambda_{g}}(\partial_{t}g_{zz})dt=-p_{0}\lambda_{g}(\partial_{t}\mu_{g})dt$ (42b)
(the minus sign $\partial_{t}g_{zz}$ in Eq. (42b) comes from the $\pi/2$-rotation in Eq. (4).) Taking the divergence, Eq. (41) becomes, in Cartesian notation
@@
$\frac{p_{0}}{2}\Delta(\dot{z})_{\beta}=\tfrac{1}{2}\partial_{\beta}(\lambda_{g}^{-1}\partial_{t}\sigma_{\alpha\alpha})-\partial_{\alpha}(\lambda_{g}^{-1}\partial_{t}\sigma_{\alpha\beta}).$ (43)
Crucially, while Eq. (43) resembles a Stokes flow, it is driven not by stress $\sigma$, but by its time-derivative $\partial_{t}\sigma$. The stress increment $(\partial_{t}\sigma_{\alpha\beta})dt$ is in the current configuration $z(t)$, and therefore not necessarily isotropic. In the new $z(t+dt)=z(t)+\dot{z}\,dt$, stress is again isotropic and can be balanced by pressure (an isotropic $\partial_{t}\sigma$ only changes the local pressure and does not induce $\dot{z}$-flow). Moreover, in contrast to a physical Stokes flow, Eq. (43) describes the “flow” of the (stress-free) reference configuration $z$.
.3.3 Physical flow
@@
We are ultimately interested in the dynamics of the physical configuration. The change of the map from reference to physical configuration is $w(z,\bar{z},t+dt)=w(z,\bar{z},t)+\dot{w}(z,\bar{z},t)\,dt$. The physical flow $\dot{w}$ is determined by the BCs and the constitutive relation $P(n)$ (see Sect. .2.4). For instance, for $B=\infty$, one obtains:
@@
Together, Eqs. (40) and (44) describe the reestablishment of force balance after an adiabatic change of stress. Flow establishes force balance, but the tensile stress remains non-zero. In this adiabatic regime, the flow rate is set by the rate of change of the tension metric, i.e. the timescale on which cells regulate contractile force generation. At this point, it would appear that a large shape change requires building up increasingly large tension anisotropy (i.e., anisotropy of tension triangles). As we will see next, however, cell rearrangements plastically lock-in past tension dynamics while resetting the built-up tension anisotropy. Macroscopically, these dynamics appear as a steady active plastic flow driven by an effective active stress.
.4 Cell rearrangement
@@
Large tissue deformations cause the cells in a tissue to rearrange, leading to plastic shape changes. In the discrete ATN model, such T1 processes correspond to edge flips that change the adjacency graph of the triangulation (Fig. 4). These flips occur when the length of a cell-cell interface vanishes $\ell_{ij}=0$ and do not manifest themselves in the tension metric. The challenge is to represent the inherently discrete T1 processes in the continuum.
Let us recall that in addition to the tension metric $\mathbf{g}$, the discrete ATN model depends on the topology of the cell tiling, i.e., the vertex adjacency matrix. To capture the adjacency information in the continuum, we introduce an “adjacency metric” $a_{\alpha\beta}(\bm{\xi})$ (defined so that adjacent cells have unit distance). It is the continuum limit of a triangulation with the same adjacency as the tension triangulation, but with all edge lengths set to unity. The metrics $\mathbf{g}(\bm{\xi})$ and $\mathbf{a}(\bm{\xi})$ define two distinct metrics on the same manifold. The Lagrangian coordinates $\bm{\xi}$ are attached to material elements (e.g. cells) and label corresponding points. The adjacency metric “geometrizes” the combinatorial information on cell-cell adjacency, and is closely related to the Koebe-Andreev-Thurston circle packing theorem (see Discussion).
Intuitively, the “discrepancy” between $\mathbf{a}$ and $\mathbf{g}$ measures the shape of tension triangles in the continuum theory, notably, the local tension anisotropy ($\mathbf{a}$ is made from equilateral “reference” triangles, $\mathbf{g}$ from potentially anisotropic tension triangles). Gaussian curvature of $\mathbf{a}$ corresponds to the net density of topological defects in a triangular lattice (i.e., 5- or 7-sided cells, Fig. 4555In an equilateral triangulation, a coordination number $\neq 6$ implies a nonzero angle deficit around a vertex, and hence non-zero Gaussian curvature concentrated at such vertices. The area element of $\mathbf{a}$ defines the local cell density, $n=n_{0}\sqrt{\det\mathbf{a}}$ (we dimensionalize $\mathbf{a}\mapsto\mathbf{a}/n_{0}$ so that $\mathbf{a}$ has units of $\mathrm{length}]^{2}$ [^6] Microscopically, the cell area in an equilateral triangulation is proportional to the coordination number. Hence, strictly speaking, $\mathbf{a}\mapsto\mathbf{a}/n_{0}$ only holds if there is no large-scale coordination number gradient..
Together, the two metrics $\mathbf{a},\,\mathbf{g}$ furnish a complete continuum description of the cell tiling and its mechanics. The ability to control tensions and topology separately is what separates the ATN model from an ordinary fluid foam, where $\mathbf{g}=\mathbf{a}$ (all tensions are equal and can be set to unity).
@@
The adjacency coordinates $\zeta$ serve as a convenient set of coordinates adapted to the adjacency of the cell network – in $\zeta$-coordinates, cells have isotropic shapes and density $n_{0}\lambda_{a}^{2}$. If interfacial tensions are perfectly isotropic, $\mathbf{a}=\mathbf{g}$ and $z=\zeta$. More generally, the Jacobian $\mathbf{D}z(\zeta,\bar{\zeta})$ of the QC map from $\zeta$ to the tension reference state $z$ defines the local tension configuration [29, 16]. Microscopically, $\mathbf{D}z(\zeta,\bar{\zeta})$ deforms equilateral “adjacency” triangles into anisotropic tension triangles. In particular, the Beltrami coefficient $\mu_{\zeta|z}=\bar{\partial}_{\zeta}z/\partial_{\zeta}z=\mu_{g}(\zeta,\bar{\zeta})$ encodes the magnitude and direction of the tension anisotropy. Overall, in $z$-coordinates, the adjacent metric simply specifies the local cell density and tension anisotropy.
By contrast, the map from $z$ to the physical configuration $w$, and its Beltrami coefficient $\mu_{z|w}=\bar{\partial}_{z}w/\partial_{z}w$, determines the elastic shear and macroscopic deviatoric stress. In fact, $\mu_{z|w}$ is proportional to deviatoric stress [cf. Eq. (20)]. The total anisotropy of the map $w(\zeta,\bar{\zeta})$ from isotropic cells to the physical configuration is given by the chain rule:
@@
where $e^{-i2\phi_{\zeta|z}}=\overline{\partial_{\zeta}z}/\partial_{\zeta}z$, and we assumed $|\mu_{\zeta|z}|,|\mu_{z|w}|\ll 1$ to make the approximation in the last step.
@@
.4.2 Kinematics of cell rearrangement
@@
$\mathbf{a}(\bm{\zeta}(t),t)\mapsto\mathbf{a}(\bm{\zeta}(t),t)+\partial_{t}\mathbf{a}(\bm{\zeta}(t),t)dt$ (49)
@@
$\partial_{t}\lambda_{a}(\zeta,\bar{\zeta},t)=0,\quad\partial_{t}\mu_{a}(\zeta,\bar{\zeta},t)=R(\zeta,\bar{\zeta},t)$ (50)
Eq. (50) can be considered a definition of the rate of topological rearrangement in the continuum.
@@
For example, consider a localized “rearrangement event” $da_{\zeta\zeta}=R\,dt=c\ell_{R}^{2}e^{-\zeta\bar{\zeta}/(2\ell_{R}^{2})}\,dt$ with spatial extent $\ell_{R}$, on an initially uniform background $\lambda_{a}=1$. The complex coefficient $c$ sets the shear orientation and magnitude. To mimic an “isolated T1” in a hexagonal lattice, one sets $\ell_{R}$ to the scale of the cell size, and $|c|\sim 1/t_{\mathrm{T1}}$, where $t_{\mathrm{T1}}$ is the timescale of the T1 process. Microscopically, such a T1 creates a 5-7 defect pair (Fig. 4, top). In the continuum, defects are localized quadrupoles in the adjacency curvature $K_{a}$. Indeed, using the St. Venant formula Eq. (13), the increment $da_{\zeta\zeta}$ leads to:
@@
For instance, for $\arg c=0$, $dK_{a}\propto(\zeta_{1}^{2}-\zeta_{2}^{2})\,dt$ (see Fig. 4, bottom).
The T1-induced shear deformation Eq. (50) requires an update to the isothermal $\zeta$-coordinates. The “topological” flow $\zeta(t)\mapsto\zeta(t+dt)=\zeta(t)+\dot{\zeta}(\zeta(t),\bar{\zeta}(t),t)dt$ obeys a Beltrami flow equation:
@@
In the updated $\zeta(t+dt)$-coordinates, $\lambda_{a}(\zeta(t+dt),\bar{\zeta}(t+dt))$ changes to accommodate the curvature increment due to $\partial_{t}\mu_{a}(\zeta(t),\bar{\zeta}(t),t)$ in the old $\zeta(t)$ coordinates.
The total time derivative of $\mathbf{g}(\bm{\zeta})$ combines two effects: intrinsic remodeling and the coordinate transformation due to the cell rearrangement flow $\zeta\mapsto\zeta+\dot{\zeta}\,dt$. The transformation law Eq. (1) implies
$\frac{d\mathbf{g}(\zeta,\bar{\zeta})}{dt}=\partial_{t}\mathbf{g}+\left[({\dot{\zeta}}\cdot\mathbf{D})\mathbf{g}-\mathbf{D}{\dot{\zeta}}\cdot\mathbf{g}-\mathbf{g}\cdot(\mathbf{D}{\dot{\zeta}})^{T}\right]$ (53)
The plastic flow of $\zeta$-coordinates appears as a continuous “reparametrization” of the tension manifold. The convective derivative of the Beltrami-coefficient $\mu_{\zeta|z}$, assuming $|\mu_{\zeta|z}|\ll 1$, is given by:
$\frac{d\mu_{\zeta|z}}{dt}$ $\approx\partial_{t}\mu_{\zeta|z}-\partial_{t}\mu_{a}+\operatorname{Re}[\bar{\dot{\zeta}}\partial_{\zeta}]\mu_{\zeta|z}$
$\qquad+2i\operatorname{Im}[\partial_{\zeta}{\dot{\zeta}}]\,\mu_{\zeta|z}$ (54)
The first two terms describe the composition of two QC transformations: by (52), $\partial_{t}\mu_{a}$ is the Beltrami coefficient of the infinitesimal map $\zeta(t)\mapsto\zeta(t+dt)$ (see Appendix F.1 for details, and Fig. 5 for an illustration of the underlying geometry). The last two terms correspond to advection and co-rotation.
To summarize the kinematic logic: T1s change the adjacency metric $\mathbf{a}$, necessitating an update of the isothermal coordinates $\zeta$. This corresponds to a reparametrization of the tension metric $\mathbf{g}$, i.e. a change of $\mu_{\zeta|z}$. This coordinate transformation leaves the tension manifold (i.e., $\mathbf{g}({\bm{\xi}})$) invariant. In our framework, cell rearrangement, thus, is a purely kinematic process and has no instantaneous effect on $\mathbf{g}$. This is best motivated by looking at physical space. Microscopically, a T1 flips an edge whose physical length $\ell_{ij}=0$, and is hence “invisible”: it has no instantaneous effect on physical shape or macroscopic stress. A specific model for the dynamic change of the tension manifold elicited by the kinematic effect of T1s on tension anisotropy will be described in the next section.
@@
To fill the kinematic relations (52)–(.4.2) with life, we need to specify a continuum law for the rate and orientation of T1s $\partial_{t}\mu_{a}$ and a model for the dynamics of $\mathbf{g}$.
@@
As a minimal, phenomenological model for topological plasticity, we propose to represent the effect of T1s by the continuous relaxation of the total anisotropy $\mu_{\zeta|w}$. The (oriented) rearrangement rate, therefore, reads
$\partial_{t}\mu_{a}(\zeta,\bar{\zeta})$ $=-\gamma_{\mathrm{T1}}\frac{\delta|\mu_{\zeta|w}|^{2}}{\delta\bar{\mu}_{a}}=\gamma_{\mathrm{T1}}\mu_{w|z}$
$\approx\gamma_{\mathrm{T1}}(\mu_{\zeta|z}+e^{-i2\phi_{\zeta|z}}\mu_{z|w})$ (55)
@@
Here, $\gamma_{\mathrm{T1}}$ is a rate coefficient (the “T1 rate”), which depends on the distribution of microscopic tension configurations, as we discuss below. The variation of $|\mu_{\zeta|w}|^{2}$ follows from the Beltrami composition formula (48) (see Appendix F.2). Equation (.4.3) implies that T1s can be driven by elastic deformations due to external forces acting on the tissue boundary, or by internal changes of tensions [29, 16]. We refer to these two scenarios as passive and active T1s. Passive T1s are driven by anisotropic elastic strain $\mu_{z|w}$ (and hence stress anisotropy $\sigma$). By contrast, active T1s can occur in the absence of macroscopic stress. Instead, they are driven by the tension anisotropy $\mu_{\zeta|z}$.
@@
Thus, Eq. (.4.3) qualitatively reproduces the mechanisms driving T1s as revealed by geometric considerations in the discrete setting [29, 15]. There, the combined tension (active) and strain (passive) anisotropy determines when a cell-cell interface reaches zero length $\ell_{ij}=0$, triggering a T1. The critical microscopic Beltrami coefficient for which $\ell_{ij}=0$ can be worked out geometrically as $|\mu_{\zeta|w}^{\triangle}|=1/2$ [15]. Here, the superscript $\triangle$ marks quantities on the single-triangle level. The Beltrami coefficients $\mu_{\zeta|z}$ and $\mu_{z|w}$ entering the continuum theory represent a macroscopic average over many tension triangles. The microscopic T1 threshold $|\mu_{\zeta|w}^{\triangle}|=1/2$ implies the existence of a yield strain for passive T1s: below a critical $|\mu_{z|w}|$, the T1 rate $\gamma_{\mathrm{T1}}$ vanishes. More generally, $\gamma_{\mathrm{T1}}$ depends on the microscopic distribution of cell-level configurations $\mu_{\zeta|w}^{\triangle}$, which must be summarized by a suitable set of local order parameters. A mean-field approach [30] then allows systematic calculation of $\gamma_{\mathrm{T1}}$ in terms of these order parameters, together with equations governing their dynamics.
.4.4 T1s relax tension and stress anisotropy

### v3

@@ Results
Eq. (42) defines a sequence of infinitesimal QC deformations – a QC flow (also called a Beltrami flow). The QC flow equation is linear and can be solved by the method of Green’s functions, even for large deformations [25] (Appendix F). Eq. (42) describes how an anisotropic change of active tension induces a corresponding shear deformation of the isothermal reference state.
For free boundaries (and under an additional condition explained below), flow of the reference state $z$ translates directly into flow of the physical configuration $w$. Eq. (42) can be brought into a more familiar form by applying $\partial_{z}$ and changing to Cartesian notation
@@
$\Delta(\dot{z})_{\beta}$ $=\partial_{\alpha}\bigl(m_{\alpha\beta}-\tfrac{1}{2}m_{\gamma\gamma}\delta_{\alpha\beta}\bigr)$ (43)
Eq. (43) resembles a Stokes flow, driven by the divergence $\mathbf{m}$. Recall, however, that $\mathbf{m}$ is not an active stress but rather the rate of change of the tension metric. In the adiabatic regime, the system follows the force-balanced configuration imposed by the internal tension configuration. The flow rate is thus set by the timescale on which cells regulate contractility, and not by the (rapid) rate $\Gamma$ of mechanical relaxation. Adiabatic flow is not stress relaxation: In the $z$-configuration, the total macroscopic stress $\bm{\sigma}^{\mathrm{tot}}=0$ throughout the adiabatic QC flow.
.3.3 Adiabatic flow of the physical configuration
@@
We are ultimately interested in the dynamics of the physical configuration $w$, which may differ from $z$ due to external or body forces. The change of the map from reference to physical configuration is $w(z,\bar{z},t+dt)=w(z,\bar{z},t)+\dot{w}(z,\bar{z},t)\,dt$. The physical flow $\dot{w}$ is determined by the BCs and the constitutive relation $P(n)$ (see Sect. .2.4). For instance, for $B=\infty$, one obtains:
$\operatorname{Re}[\partial_{z}\dot{w}]$ $=0$ (44a)
$\partial_{z}\bar{\partial}_{z}\operatorname{Im}[\partial_{z}(\dot{z}-\dot{w})]$ $=0$ (44b)
@@
Together, Eqs. (42) and (44) describe the reestablishment of force balance after a change of stress. At this point, it would appear that a large shape change requires building up increasing tension anisotropy (i.e., tension triangles anisotropy). As we will see next, however, cell rearrangements plastically lock in past tension dynamics while resetting the built-up tension anisotropy. Macroscopically, these dynamics appear as a steady plastic flow.
.4 Topological rearrangement
@@
Large tissue deformations cause cells to rearrange, leading to plastic shape changes. In the discrete model, such T1 processes correspond to edge flips that change the adjacency graph of the triangulation (Fig. 4). (“T1” stands for “topological 1”; in the following, we use “topology” to refer to the adjacency graph and not the gross topology of the tension manifold, which we assume to be that of a disk). These flips occur when the length of a cell-cell interface vanishes $\ell_{ij}=0$ and do not manifest themselves in the tension metric. The challenge is to represent the inherently discrete T1 processes in the continuum. We now introduce a mathematical framework to do so and return to the physical consequences of T1s for plastic flow in Sects. .4.3-.4.4.
Let us recall that in addition to the tension metric $\mathbf{g}$, the discrete ATN model depends on the topology of the cell tiling, i.e., the vertex adjacency matrix. To capture the adjacency information in the continuum, we introduce an “adjacency metric” $a_{\alpha\beta}(\bm{\xi})$ defined so that adjacent cells have unit distance. It is the continuum limit of a triangulation with the same adjacency as the tension triangulation, but with all edge lengths set to unity. Thus $\mathbf{g}(\bm{\xi})$ and $\mathbf{a}(\bm{\xi})$ define two distinct metrics on the same manifold. The Lagrangian coordinates $\bm{\xi}$ are attached to material elements (e.g. cells) and label corresponding points. The adjacency metric “geometrizes” the topological information on cell-cell adjacency, and is closely related to the Koebe-Andreev-Thurston circle packing theorem (see Discussion).
Intuitively, the “discrepancy” between $\mathbf{a}$ and $\mathbf{g}$ measures the shape of tension triangles in the continuum theory, notably, the local tension anisotropy ($\mathbf{a}$ is made from equilateral “adjacency” triangles, $\mathbf{g}$ from potentially anisotropic tension triangles). Gaussian curvature of $\mathbf{a}$ corresponds to the net density of topological defects in a triangular lattice (i.e., 5- or 7-sided cells, Fig. 4666In an equilateral triangulation, a coordination number $\neq 6$ implies a nonzero angle deficit around a vertex, and hence non-zero Gaussian curvature concentrated at such vertices. The area element of $\mathbf{a}$ defines the local cell density, $n=n_{0}\sqrt{\det\mathbf{a}}$ (we dimensionalize $\mathbf{a}\mapsto\mathbf{a}/n_{0}$ so that $\mathbf{a}$ has units of $\mathrm{length}]^{2}$ [^7] Microscopically, the cell area in an equilateral triangulation is proportional to the coordination number. Hence, strictly speaking, $\mathbf{a}\mapsto\mathbf{a}/n_{0}$ only holds if there is no large-scale coordination number gradient..
Together, $\mathbf{a}$ and $\mathbf{g}$ furnish a complete continuum description of the cell tiling and its mechanics. The ability to control tensions and topology separately is what separates the ATN model from an ordinary fluid foam, where $\mathbf{g}=\mathbf{a}$ (all tensions are equal and can be set to unity).
@@
The adjacency coordinates $\zeta$ serve as a convenient set of coordinates adapted to the adjacency of the cell network – in $\zeta$-coordinates, cells have isotropic shapes and density $n_{0}\lambda_{a}^{2}$. If interfacial tensions are perfectly isotropic, $\mathbf{a}=\mathbf{g}$ and $z=\zeta$. More generally, the Jacobian $\mathbf{D}z(\zeta,\bar{\zeta})$ of the QC map from $\zeta$ to the tension reference state $z$ defines the local tension configuration [38, 19]. Microscopically, $\mathbf{D}z(\zeta,\bar{\zeta})$ deforms equilateral “adjacency” triangles into anisotropic tension triangles. In particular, the Beltrami coefficient $\mu_{\zeta|z}=\bar{\partial}_{\zeta}z/\partial_{\zeta}z=\mu_{g}(\zeta,\bar{\zeta})$ encodes the magnitude and direction of the tension anisotropy. Overall, in $z$-coordinates, the adjacent metric specifies the local cell density and the tension anisotropy.
By contrast, the map from $z$ to the physical configuration $w$, and its Beltrami coefficient $\mu_{z|w}=\bar{\partial}_{z}w/\partial_{z}w$, determines the elastic shear and macroscopic deviatoric stress. In fact, $\mu_{z|w}$ is proportional to deviatoric stress [cf. Eq. (20)]. The total anisotropy of the map $w(\zeta,\bar{\zeta})$ from isotropic adjacency triangles to the physical configuration is given by the chain rule:
@@
where $e^{-i2\phi_{\zeta|z}}=\overline{\partial_{\zeta}z}/\partial_{\zeta}z$, and we assumed $|\mu_{\zeta|z}|,|\mu_{z|w}|\ll 1$ in the last step.
@@
.4.2 Kinematics of cell rearrangement: Flip flow
@@
$\mathbf{a}\bigl(\bm{\zeta}(t),t\bigr)\mapsto\mathbf{a}\bigl(\bm{\zeta}(t),t\bigr)+\partial_{t}\mathbf{a}\bigl(\bm{\zeta}(t),t\bigr)dt$ (49)
@@
$\partial_{t}\lambda_{a}(\zeta,\bar{\zeta},t)=0,\quad\partial_{t}\mu_{a}(\zeta,\bar{\zeta},t):=R(\zeta,\bar{\zeta},t)$ (50)
In the present framework, Eq. (50) is the definition of the rate of topological rearrangement in the continuum. We introduced $R$ as a complex Beltrami coefficient; it can also be represented by a (trace-free) Cartesian tensor $\mathbf{R}$ analogous to $\mathbf{m}$ in Eq. (36).
@@
For example, consider a localized “rearrangement event” $da_{\zeta\zeta}=\lambda_{a}^{2}R\,dt=c\ell_{R}^{2}e^{-\zeta\bar{\zeta}/(2\ell_{R}^{2})}\,dt$ with spatial extent $\ell_{R}$, on an initially uniform background $\lambda_{a}=1$. The complex coefficient $c$ sets the shear orientation and magnitude. To mimic an “isolated T1” in a hexagonal lattice, one sets $\ell_{R}$ to the scale of the cell size, and $|c|\sim 1/t_{\mathrm{T1}}$, where $t_{\mathrm{T1}}$ is the timescale of the T1 process. Microscopically, such a T1 creates a 5-7 defect pair (Fig. 4, top). In the continuum, defects are localized quadrupoles in the adjacency curvature $K_{a}$. Indeed, using the St. Venant formula Eq. (13), the increment $da_{\zeta\zeta}$ leads to:
@@
For $\arg c=0$, $dK_{a}\propto(\zeta_{1}^{2}-\zeta_{2}^{2})\,dt$ (Fig. 4, bottom).
The T1-induced shear deformation Eq. (50) requires an update to the isothermal $\zeta$-coordinates, $\zeta(t)\mapsto\zeta(t+dt)=\zeta(t)+\dot{\zeta}(\zeta(t),\bar{\zeta}(t),t)dt$. We refer to $\dot{\zeta}(t)$ as the flip flow, since it corresponds to edge flips in the discrete triangulation. The flip flow is quasi-conformal and obeys:
@@
The time derivative of $\mathbf{g}(\bm{\zeta})$ combines intrinsic active tension dynamics $m$ and cell rearrangement $R$. The former leads to reference state flow $z\mapsto z+\dot{z}\,dt$, and the latter to the coordinate transformation $\zeta\mapsto\zeta+\dot{\zeta}\,dt$. The flip flow thus enters as a continuous “reparametrization” of the tension manifold. The dynamics of the Beltrami-coefficient $\mu_{\zeta|z}$, assuming $|\mu_{\zeta|z}|\ll 1$, is given by:
$\frac{d}{dt}\mu_{\zeta|z}$- $\left(\mathrm{Re}[\dot{\zeta}\partial_{\zeta}]\mu_{\zeta|z}+i2\mu_{z}\mathrm{Im}[\partial_{\zeta}{\dot{\zeta}}]\right)\approx$
$\quad\partial_{t}\mu_{g}(\zeta,\bar{\zeta})-\partial_{t}\mu_{a}(\zeta,\bar{\zeta})$ (53)
The bracketed terms on the LHS represent advection and co-rotation and will be suppressed in the following. The RHS arises from the QC transformations of $z$ and $\zeta$ (see Appendix G.1 for the derivation; Fig. 8 illustrates the underlying geometry.).
In summary, T1s change the adjacency metric $\mathbf{a}$, which we represent via the flip flow $\dot{\zeta}$. The flip flow is a reparameterization that changes the tension anisotropy, but has no instantaneous effect on the tension metric $\mathbf{g}({\bm{\xi}})$ in fixed, Lagrangian coordinates. In our framework, cell rearrangement is thus a purely kinematic process. This is best motivated by looking at physical space. Microscopically, a T1 flips an edge whose physical length $\ell_{ij}=0$, and is hence “invisible”: it has no instantaneous effect on physical shape or macroscopic stress. We next describe a specific model for the dynamic change of the tension manifold elicited by the kinematic effect of T1s on tension anisotropy.
@@
To fill the kinematic relations (52)–(.4.2) with life, we need to specify a continuum law for the rate and orientation of T1s $R$ and a prescription for the dynamics of $\mathbf{g}$ caused by rearrangements.
@@
As a minimal, phenomenological model for topological plasticity, we propose to represent T1s by continuous relaxation of the total anisotropy $\mu_{\zeta|w}$. The (oriented) rearrangement rate, therefore, reads
$R(\zeta,\bar{\zeta})$ $=-\gamma_{\mathrm{T1}}\frac{\delta|\mu_{\zeta|w}|^{2}}{\delta\bar{\mu}_{a}}=\gamma_{\mathrm{T1}}\mu_{\zeta|w}$ (54)
@@
Here, $\gamma_{\mathrm{T1}}$ is a rate coefficient (the “T1 rate”), which depends on the distribution of microscopic tension configurations, as we discuss below. The variation of $|\mu_{\zeta|w}|^{2}$ follows from the Beltrami composition formula (48) (see Appendix G.2). The total anisotropy decomposes into the tension anisotropy and the stress anisotropy
$\mu_{\zeta|w}\approx\mu_{\zeta|z}+e^{-i2\phi_{\zeta|z}}\mu_{z|w}.$ (55)
T1s can thus be driven by internal tension changes or by elastic deformations due to external forces [38, 19]. Accordingly, we define active and passive T1 rates
$R_{\mathrm{act}}$ $=\gamma_{\mathrm{T1}}\mu_{\zeta|z}$ (56a)
$R_{\mathrm{pass}}$ $=\gamma_{\mathrm{T1}}\mu_{z|w}$ (56b)
so that $R=R_{\mathrm{act}}+e^{-i2\phi_{\zeta|z}}R_{\mathrm{pass}}$ Passive T1s are driven by anisotropic elastic strain $\mu_{z|w}$ (and hence stress anisotropy $\sigma_{ww}$). By contrast, active T1s can occur in the absence of macroscopic stress. Instead, they are driven by the tension anisotropy $\mu_{\zeta|z}$.
@@
Thus, Eq. (54) qualitatively reproduces the mechanisms driving T1s on the cell level [38, 18]. There, the combined tension (active) and strain (passive) anisotropy determines when a cell-cell interface reaches zero length $\ell_{ij}=0$, triggering a T1. The critical microscopic Beltrami coefficient for which $\ell_{ij}=0$ can be worked out geometrically as $|\mu_{\zeta|w}^{\triangle}|=1/2$ [18]. Here, the superscript $\triangle$ marks quantities on the single-triangle level. The Beltrami coefficients $\mu_{\zeta|z}$ and $\mu_{z|w}$ entering the continuum theory represent a macroscopic average over many tension triangles. The microscopic T1 threshold $|\mu_{\zeta|w}^{\triangle}|=1/2$ implies the existence of a yield strain for passive T1s: below a critical $|\mu_{z|w}|$, the T1 rate $\gamma_{\mathrm{T1}}$ vanishes. More generally, $\gamma_{\mathrm{T1}}$ depends on the microscopic distribution of cell-level configurations $\mu_{\zeta|w}^{\triangle}$, which must be summarized by a suitable set of local order parameters. A mean-field approach [39] allows systematic calculation of $\gamma_{\mathrm{T1}}$ in terms of these order parameters, together with equations governing their dynamics.
.4.4 T1s relax tension- and stress anisotropy